APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Sunday January 31, 2010 (MJD 55228) ---=== OBSERVING TEAM ===--- Audrey Simmons Elena Malanushenko Kaike Pan (afternoon) Support: Steph Snedden Paul Harding ---=== OBSERVING PLAN ===--- MARVELS Science! ---=== OBSERVING SUMMARY ===--- We did not open as the humidity was high and it was cloudy. Paul worked with guider problems. Spectro dewar needs to be replaced in the morning. Marvels ch-8 CCD has a little bit little temperature but this created warnings. Marvels team needs to check it. Spectro chilly gauge shows zero again and needs to be checked. ---=== OBSERVING LOG ===--- Afternoon: ---------- All systems and telescopes were checked, no problem to report. A pair of BOSS bias and dark were taken, but SOS3 does not show reduced results. So we stopped running processes (4 of them), and re-start the sos. 4 processed are running. 23:30Z Marvels Thar lamps on. High humidity prevents afternoon ventilation. Paul is working on catridge 3 issue and B-side only plate issue. Night: ----- There are clouds and >86% humidity with ~2C diff. We delay with the openings. Cart 7 is mounted. I ack warns alerts: Warn boss.SP1B2IPosTrim2 Reported by camCheck Warn boss.SP1R0INegTrim1 Reported by camCheck Warn boss.SP1R0IPlusTrim1 Reported by camCheck Warn boss.SP1R0IPlusTrim2 Reported by camCheck Warn boss.SP1R0TZero Reported by camCheck Warn boss.SP2R0VRD2 Reported by camCheck and disabled alert: Critical imager.heartbeat APO.elenam Marvels CCD a little bit out of temperature range: Error: ALARM> ch-8 : CCD house = 28.051 > 28.05 C Spectro chilly gauge show zero again similar as it was recently. Somebody need to check it. Paul was working with guider simulation for whole night. Several times cartridges with side B were unavailble, we made them active. The humidity was high for whole night, we did not open. 11:40 telescope was off, 11:45 endNight is completed. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 Boss (cartridge is not important):: 1 23:05 b1-00107967 4 3413 bia Out 1.3 19864 20469 20279 2 23:09 b1-00107968 4 3413 bia Out 1.3 19864 20469 20279 3 23:13 b1-00107969 4 3413 dar Out 21.3 19864 20469 20279 ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 23:23Z 31 27 10 276 (W) 304 - 23:53Z 31 27 11 283 (WNW) 339 - 00:23Z 30 24 11 281 (W) 282 - 01:06Z 28 24 6 268 (W) 287 - 01:36Z 27 25 6 275 (W) 298 - 02:06Z 27 25 11 304 (NW) 306 - 02:37Z 27 25 13 313 (NW) 295 - 03:07Z 27 26 10 311 (NW) 322 - 03:37Z 28 26 8 289 (WNW) 302 - 04:10Z 28 26 13 297 (WNW) 342 - 04:41Z 28 27 12 284 (WNW) 378 - 05:11Z 28 27 12 287 (WNW) 394 - 05:45Z 28 27 6 294 (WNW) 368 - 06:19Z 28 25 12 272 (W) 358 - 06:53Z 27 26 6 266 (W) 370 - 07:23Z 27 27 6 272 (W) 3188 - 07:57Z 27 27 9 281 (W) 112478 - 08:31Z 26 26 9 312 (NW) 23994 - 09:05Z 26 25 9 295 (WNW) 293 - 09:35Z 27 24 13 304 (NW) 272 - 10:09Z 27 24 11 316 (NW) 293 - 10:39Z 26 24 16 323 (NW) 275 - 11:13Z 26 23 8 329 (NNW) 294 - 11:43Z 26 23 8 346 (NNW) 314 - ---=== TELESCOPE STATUS ===--- Did not open last night. No interlocks bypassed. Status at: 11:40Z Telescope stowed at: 30,121,00 mount Instrument mounted: Cartridge 7 Counterweights at: 240 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: Spectro 43 lbs, 20 psi Imager lbs, psi (IMAGER IS WARM) ---=== SOFTWARE USED ===--- IOP/SOP: v4_40_0 Watcher: v2_36_0 MCP: v6_6_3 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: TCC 2.11.0 2009-06-24 sdssProcedures: SoS: plate-mapper: STUI 0.4a7 sdss3@hub25m[640]$ getVersions guider: v1_0_16 current sop: v1_0_4 current gcameraICC: v1_0_5 current alerts: v1_0_3 current platedB: rhl current ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -0.7010 -0.6940 Axial B -0.0240 -0.0410 Axial C -0.3770 -0.3710 Trans D 11.6110 11.6170 Lateral E 10.3759 10.3505 Lateral F 11.9507 11.9253 GALILS Commanded: 9900. 400. 5850. 1200. 2950. 2750. Actual: 9900. 400. 5850. 1200. 2950. 2750. SETMIR VALUES PriDesOrient: 0.00 -3.30 5.10 278.08 178.06 PriOrient: 0.00 -3.04 5.26 277.70 178.07 SECONDARY: ---------- Focus: 0.00 Air Temp.: -1.2 Alt.: 29.999997 MIGS TONIGHT NOMINAL Axial A 18.6180 18.6150 Axial B 15.6280 15.6120 Axial C 13.8250 13.8170 Trans D -11.7750 -11.7780 GALILS Commanded: 1598918. 1565110. 1586824. -7600. -12400. Actual: 1597700. 1566200. 1588100. -7600. -12400. SETMIR VALUES SecDesOrient: 1257.00 -4.00 -9.00 10.00 234.42 SecOrient: 1256.69 -3.68 -9.77 9.68 234.48 ---=== PROBLEMS IN DETAIL ===---